A tuition at Rutgers University was $38,148 in 1995. In 2002, the tuition was $42,727.

a)Assuming tuition is increasing at an exponential rate, write a function f(y) representing the cost of tuition y years since 1995.
b)What is the growth rate per year?
c) What is the growth rate per decade?
d) write function f(d) representing the cost of tuition d decades since 1995.
e) if this trend were to continue, what should the tuition be in the year 2020?
1
Expert's answer
2020-03-30T12:38:47-0400

a) Let "y" be the number of years since 1995 and let "f(y)" be the cost of tuition in dollars. 

Exponential growth function


"f(y)=f(0)e^{ky}"

Given that "f(0)=\$38148." Then


"f(y)=38148e^{ky}"

b) A tuition at Rutgers University was $42,727 in 2002.


"2002-1995=7"

"f(7)=38148e^{7k}=42727"

Solve for "k"


"e^{7k}={42727 \over 38148}"

"\ln(e^{7k})=\ln\big({42727 \over 38148}\big)"

"7k=\ln\big({42727 \over 38148}\big)"

"k={1 \over 7}\ln\big({42727 \over 38148}\big)"

"k\approx0.016194"

"f(y)=38148e^{0.016194y}=38148(1.016326)^y=38148(1+0.016326)^y"

The growth rate per year between 1995 and 2002


"r=0.016326"

c) Let "d" be the number of decades since 1995 and let "f(d)" be the cost of tuition in dollars. 


"f(d)=38148e^{md}"

A tuition at Rutgers University was $42,727 in 2002


"f({7 \over 10})=38148e^{m({7 \over 10})}=42727"

Solve for "m"


"e^{0.7m}={42727 \over 38148}"

"\ln(e^{0.7m})=\ln\big({42727 \over 38148}\big)"

"0.7m=\ln\big({42727 \over 38148}\big)"

"m={10 \over 7}\ln\big({42727 \over 38148}\big)"

"m\approx0.161940"

"f(d)=38148e^{0.16194d}=38148(1.175789)^d=38148(1+0.175789)^d"

The growth rate per decade between 1995 and 2002


"q=0.175789"

d)


"f(d)=38148e^{0.16194d}=38148(1.175789)^d=38148(1+0.175789)^d"

e)


"2020-1995=25"

"f(25)=38148e^{0.016194(25)}=\$57186.81"

A tuition at Rutgers University should be "\$57186.81" in 2020.